Apr 24, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and SeeVogh session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  • BCAL: Check max ratings for MPPCs (Elton).

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting May 12-14
  3. Action Items
  4. FCAL
  5. BCAL
  6. Reconstruction and Simulation
  7. TDR update
  8. Calibration
  9. Any other business


Attendees: Elton, John, Yi, Andrei, Will, David (JLab); Shaun (UofR); Christina, George, Pavols (Athens).

  1. Announcements
  2. Collaboration Meeting May 12-14
    • John will give an update on the Fcal and Elton will give an update on the Bcal.
  3. Action Items
  4. Monitoring (Athens)
    • Christina reported that George Vasileiadis visited the university during his Easter break. He is likely to be stationed in Athens for some time and can stay in touch during that time.
  5. FCAL HDFCAL log book
    1. FADC problems
      • Working with Hai and Brian
      • Multiple problems encountered with Flashes. Five Fadcs replaced. Only one actually had a hardware problem.
      • New firmware update fixed all problems on the north side (likely on south also, but not checked yet)
      • Plots show that all signals from the LED pulsing system are within about 20 ns of each other in the Fadc. Delays are systematic across the face of the Fcal due to propagation of LED light through the acrylic
    2. Next steps
      • Recheck all channels, including south side
      • Gain balancing is next. Plan to use LED light to smooth out response between neighbors. Response currently varies by factors of 2-3.
      • Possibly use cosmics to check response to muons
      • The goal is to minimize the differences in response so that calibrations using beam pi0s will converge.
      • Use TOF counters to trigger events and see if this can help gain matching
    3. Response to muons
      • Test at IU to determine the response of calorimeter to cosmic muons has been setup, allowing the mini-Fcal to be pointed toward the sky.
      • Hardware is in place, but DAQ is not ready
      • Student working on project will be at JLab this summer
  6. BCAL HDBCAL log book
    1. Update
      • Upstream and downstream modules are operational. One channel 42D A8 has a large baseline offset and has been removed for Chris and Nick to fix.
      • Cooling system is operational and the interlocks working
      • DAQ is being checked on the north side (the trigger group is using the south side)
      • Pedestals have been aligned on 3 out of 4 crates
      • Recent problem with readout gives different number of output words for different crates with the same number of cards
      • (John): There are several reasons that can cause this problem. Check that the DAC values are correct and none missing. Check that the number of fADCs being readout is correct. Check that all slots are being readout
    2. Cosmic-ray setup (Andrei)
      • Identified two scintillators from the FDC cosmic setup that we could use. The counters are being installed under the magnet perpendicular to the beam. This will allow us to trigger on cosmic rays that cover a short distance along the beamline (z). The two counters are 120x20cm2 and match the area of the Qweak counter on top of the magnet.
      • The counters were mounted on separate backing structures for transport and support under the magnet.
      • Electronics and cables have been identified to hook up the scintillators, but they have not yet been installed. Most of this should be done by tomorrow.
    3. Update on ESB cosmic testing (Shaun) Media:cosmics.pdf
      • Analysis of all 49 modules has been completed
      • There is a double peaking in the distribution of average response for different modules, which is not quite understood. It was suggested that some of this could be due to temperature excursions, but this does not explain all the data.
      • The relative gains within one module seem to be very consistent, i.e. between upstream/downstream and layers 4,3,2 vs 1.
      • Draft of thesis has been submitted to committee. A final version including the analysis of all 49 modules and feedback from his committee will be produced in about a month. Thesis defense is mid May.
  7. Reconstruction and Simulation
    • Next week's meeting will be a review of status introducing Teagan to the project
    • Will Levine will give a survey of where things stand. He has recently committed changes to svn that remove 'old BCAL simulation code'.
    • Teagan will take up the project this summer
  8. TDR update
    • Zisis produced a BCAL draft, which has been checked in
    • Elton is revising the draft and will work with Zisis to have a draft that can be circulated more widely by the collaboration meeting.
    • No update on the FCAL draft.
  9. Calibration
  10. Any other business